Kate is three years less than twice as old as her brother, Brett. Next year Brett's age will be three fourths Kate's age. How old is each now?

Answer by mananth(16949)

let the brother's age be xyears
twice the age = 2x
Kate will be 2x-3 years.
.
Nest year their ages will be
Brett = x+1
kate = 2x-3 +1
.
x+1 = 3/4 ( 2x-2)
4(x+1) = 3(2x-2)
4x+4=6x-6
2x=10
x=5 Brett's age is 5 years
Kate's age will be x-3 = 5-3 =2 years